OpenStudy (anonymous):

The lengths of the legs of a right triangle are 15 cm and 20 cm. What is the perimeter of the triangle? Note: You will need to use the Pythagorean theorem to find the hypotenuse. A. 35 cm B. 25 cm C. 60 cm D. 300 cm E. 20 cm

OpenStudy (mertsj):

did you find the hypotenuse yet?

OpenStudy (anonymous):

the hypotenuse is 25 and its perimeter is 60

OpenStudy (anonymous):

c^2=a^2+b^2 <= hypotenuse a+b+c=Perimeter